OpenStudy (onepieceftw):

Find sin θ if cot θ = - 2 and cos θ < 0. Do I use the reciprocal identity cot = 1/tan θ? Or do I use the quotient identity cot θ = cosθ/sinθ?

OpenStudy (anonymous):

reciprocal

OpenStudy (superteenblue):

hi

OpenStudy (superteenblue):

hello

OpenStudy (superteenblue):

hi

OpenStudy (onepieceftw):

Hi. Can you help?

OpenStudy (superteenblue):

no

OpenStudy (superteenblue):

need help

OpenStudy (onepieceftw):

Would you not clutter other people's question with your own? It's very annoying.

OpenStudy (superteenblue):

I hate my little brother he went on my computer and talk to people about nasty stuff yesterday

OpenStudy (superteenblue):

ok

OpenStudy (welshfella):

draw a diagram 

OpenStudy (welshfella):

now use pythagoras to find measure of the hypotenuse .

OpenStudy (onepieceftw):

\[\sqrt{5}\]

OpenStudy (welshfella):

right so sin theta = ?

OpenStudy (onepieceftw):

\[\frac{ 1 }{ \sqrt{5} }\]

OpenStudy (welshfella):

yep

OpenStudy (onepieceftw):

How do I then rationalize the denominator?

OpenStudy (welshfella):

oh ok multiply topand bottom by sqrt5

OpenStudy (onepieceftw):

Okay, awesome. Thank bud

OpenStudy (welshfella):

yw